# PersonaBooth: Personalized Text-to-Motion Generation

### v1

@@ 1 Introduction
Another challenge is extracting a consistent persona across different atomic motions, which manifests in diverse ways depending on the action content. For example, the elegance of a ballerina is expressed in her feet as she walks and in her hands as she waves. MST methods have faced similar challenges when transferring styles between different content. Approaches like those of Park et al. [20], Jang et al. [9], and MoMo [25] leave this as a limitation. In contrast, some works have attempted to solve this problem. For instance, MoST [12] introduces a style disentanglement loss, and MCM-LDM [32] explicitly uses the target motion’s trajectory. However, MoST’s loss function requires running the model twice for two different input motions, which is computationally expensive, especially for diffusion models with numerous steps. Additionally, since our goal is to generate motion from text rather than from target motions, the method used by MCM-LDM is not applicable. Therefore, we introduce a novel contrastive learning-based loss specifically designed for text-to-motion diffusion, called persona cohesion loss. This loss facilitates cohesion across 으로 사사motions with different content but the same persona.
@@ 3.2 Textual and Visual Adaptation
The structure of the Personalized Text Encoder ($\mathcal{X}$) is shown in Fig. 3 (a). Given an input sentence like “A person is dancing” ($T_{in}$), we introduce a personalized sentence “[P*] person is dancing” ($\tilde{T}_{in}$), where the word [P*] precedes the subject. Subsequently, the vector $P^{*}$, derived from $\mathcal{E}$, replaces the token embedding corresponding to [P*]. Then, the modified sentence embedding, $\mathcal{X}_{clip}(\tilde{T}_{in})$, is adaptively merged with the original sentence embedding, $\mathcal{X}_{clip}(T_{in})$. The descriptions in HumanML3D are generally structured with subjects like ‘A Person,’ or ‘Someone’ and lack adjectives to characterize the subject. Thus, a smoother adaptation is necessary to enable the pretrained model to handle variations in sentence structure effectively. Specifically, personalized text feature $T^{*}$ is derived as
@@ 4 PerMo Dataset
Figure 5: Example of the ablation study. The input motions are from the ‘Uppity’ of Actor 1. The input prompt is “A person walks in a circle.” In (a) and (b), only $M_{1}$ is provided for the input, while both $M_{1}$ and $M_{2}$ are provided for (c) and (d). $L_{pc}$ encourages the generated motion to closely follow the prompt, while CAF prevents the motion from blending. We set $k=1$ for CAF
@@ 5.1 Ablation Study
Figure 6: Example reflecting the different personas of Actors 1-5 for the same ‘Childish’ category. The input prompt is “A person is running and bending to pick something.” Red arrows highlight the frames that clearly show distinct movements. The input motion of each actor can be referenced in Fig. 4 (b)
@@
Figure 7: Qualitative comparison on PerMo dataset. The input prompt is “A person hops forward and turns in the air.”
Figure 8: Qualitative comparison on 100Style dataset. The input prompt is “A person walks backward and sits down on the chair.”
@@ 5.2.1 PerMo Dataset
Fig. 8 shows qualitative results. MoMo [25] notes that their generated motions often follow the input motion’s direction instead of the target’s, which they attempted to copy the root rotation from the input motion (shown in (a) before and (b) after adjustment). The prompts for “turns in the air” were not met in (a), (b), or (c). Particularly, (c) does not reflect the style of ballerina. In contrast, PersonaBooth accurately generates the motion by first performing a hop that reflects the style, followed by a precise turn in the air. Example videos are provided in the supplementary material.
@@ 6 Conclusion
Limitations and Future Works. The PerMo contains relatively short motion sequences, which sometimes lead to stationary poses at the end of generated motions of the same length. However, this issue does not occur when prompted with longer actions, suggesting that automatic adjustment of motion length remains a task for future work. In addition, we plan to investigate approaches for applying CAF to each sequential action within a prompt individually. For instance, applying CAF separately to ‘run’ and ‘jump’ in the prompt “A person runs and then jumps.”
@@ A.2 Persona and Content Categories
Table A: Style categories included in the PerMo dataset
@@
Table B: Content categories included in the PerMo dataset
@@ A.4 Data Format and Post-Processing
Marker [style]_[content]_[actor]_[take].bvh 41 markers for each frame
Skeleton [style]_[content]_[actor]_[take].c3d 20 bones for each frame
@@
Table C: Description of data format and naming conventions
@@ A.6 Text Description
Examples of text descriptions included in the PerMo dataset are presented in Fig. G. To generate diverse descriptions, we first provided ChatGPT with detailed explanations of the motions. For example, for the ‘Hop’ motion, we used a description like: “In the motion, one person hops forward on one leg. The person hops several times.” We then instructed ChatGPT to create 20–30 variations of sentences, ranging from short and simple high-level descriptions to long and detailed low-level ones. As a result, we obtained descriptions of the ‘Hop’ motion at various levels of detail, as shown in Fig. G (a). Additionally, we represented the placeholder word [P*] within the sentences using sks [30]. When extracting sentence features, the word embedding corresponding to the sks index is replaced with $P^{*}$.
@@ C Ablation Study on the Number of Inputs
Figure I: numinput
As $|M_{i}|$ increases, all metrics demonstrate an increasing trend. This can be attributed to the higher likelihood of encountering motions that are more contextually aligned with the prompt as the diversity of input motions grows.
@@ D Persona Recognition Accuracy Details
For PerMo, PRA aims to classify 170 personas. To reduce the burden on a single classifier handling too many classes, we used nine separate classifiers, each responsible for a sub-parent category. Table D shows these sub-parent categories for PRA. For instance, the classifier for the ”Age” category handles 20 personas, with each style containing five different personas.
Each classifier is a 2-block transformer-based model, pretrained separately for each sub-parent category. The classifiers are trained using ground truth motions in PerMo paired with persona labels. These classifiers evaluate the generated motions, and their classification accuracy determines the PRA. Note that all 170 personas are combined into a single dataset during training.
For 100Style, Style Recognition Accuracy (SRA) is evaluated across 40 style categories, excluding content-oriented ones. Only one classifier is used for 100Style.

### v3

@@ 1 Introduction
Another challenge is extracting a consistent persona across different atomic motions, which manifests in diverse ways depending on the action content. For example, the elegance of a ballerina is expressed in her feet as she walks and in her hands as she waves. MST methods have faced similar challenges in disrupting the content of the target motion when transferring styles from the source motion with different content. Approaches like those of Park et al. [20], Jang et al. [9], and MoMo [25] cite this as a limitation. In contrast, MoST [12] addresses the issue with a style disentanglement loss, while MCM-LDM [32] uses the target motion’s trajectory to preserve content. However, MoST’s loss function requires running the model twice for two different input motions, which is computationally expensive, especially for diffusion models with numerous steps. Additionally, since our goal is to generate motion from text rather than from target motions, the method used by MCM-LDM is not applicable. Therefore, we introduce a novel contrastive learning-based loss specifically designed for text-to-motion diffusion, called persona cohesion loss. This loss facilitates cohesion across motion features with different content but the same persona.
@@ 3.2 Textual and Visual Adaptation
The structure of the Personalized Text Encoder ($\mathcal{X}$) is shown in Fig. 3 (a). Given an input sentence like “A person is dancing" ($T_{in}$), we introduce a personalized sentence “[P*] person is dancing" ($\tilde{T}_{in}$), where the word [P*] precedes the subject. Subsequently, the vector $P^{*}$, derived from $\mathcal{E}$, replaces the token embedding corresponding to [P*]. Then, the modified sentence embedding, $\mathcal{X}_{clip}(\tilde{T}_{in})$, is adaptively merged with the original sentence embedding, $\mathcal{X}_{clip}(T_{in})$. The descriptions in HumanML3D are generally structured with subjects like ‘A Person,’ or ‘Someone’ and lack adjectives to characterize the subject. Thus, a smoother adaptation is necessary to enable the pretrained model to handle variations in sentence structure effectively. Specifically, personalized text feature $T^{*}$ is derived as
@@ 4 PerMo Dataset
Figure 5: Example of the ablation study. The input motions are from the ‘Uppity’ of Actor 1. The input prompt is “A person walks in a circle." In (a) and (b), only $M_{1}$ is provided for the input, while both $M_{1}$ and $M_{2}$ are provided for (c) and (d). $L_{pc}$ encourages the generated motion to closely follow the prompt, while CAF prevents the motion from blending. We set $k=1$ for CAF
@@ 5.1 Ablation Study
Figure 6: Example reflecting the different personas of Actors 1-5 for the same ‘Childish’ category. The input prompt is “A person is running and bending to pick something." Red arrows highlight the frames that clearly show distinct movements. The input motion of each actor can be referenced in Fig. 4 (b)
@@
Figure 7: Qualitative comparison on PerMo dataset. The input prompt is “A person hops forward and turns in the air."
Figure 8: Qualitative comparison on 100Style dataset. The input prompt is “A person walks backward and sits down on the chair."
@@ 5.2.1 PerMo Dataset
Fig. 8 shows qualitative results. MoMo [25] notes that their generated motions often follow the input motion’s direction instead of the target’s, which they attempted to copy the root rotation from the input motion (shown in (a) before and (b) after adjustment). The prompts for “turns in the air" were not met in (a), (b), or (c). Particularly, (c) does not reflect the style of ballerina. In contrast, PersonaBooth accurately generates the motion by first performing a hop that reflects the style, followed by a precise turn in the air. Example videos are provided in the supplementary material.
@@ 6 Conclusion
Limitations and Future Works. The PerMo contains relatively short motion sequences, which sometimes lead to stationary poses at the end of generated motions of the same length. However, this issue does not occur when prompted with longer actions, suggesting that automatic adjustment of motion length remains a task for future work. In addition, we plan to investigate approaches for applying CAF to each sequential action within a prompt individually. For instance, applying CAF separately to ‘run’ and ‘jump’ in the prompt “A person runs and then jumps."
@@ A.2 Persona and Content Categories
Table A: Style categories included in the PerMo dataset
@@
Table B: Content categories included in the PerMo dataset
@@ A.4 Data Format and Post-Processing
Table C: Description of data format and naming conventions
@@
Marker [style]_[content]_[actor]_[take].c3d 41 markers for each frame
Skeleton [style]_[content]_[actor]_[take].bvh 20 bones for each frame
@@ A.6 Text Description
Examples of text descriptions included in the PerMo dataset are presented in Fig. G. To generate diverse descriptions, we first provided ChatGPT with detailed explanations of the motions. For example, for the ‘Hop’ motion, we used a description like: “In the motion, one person hops forward on one leg. The person hops several times." We then instructed ChatGPT to create 20–30 variations of sentences, ranging from short and simple high-level descriptions to long and detailed low-level ones. As a result, we obtained descriptions of the ‘Hop’ motion at various levels of detail, as shown in Fig. G (a). Additionally, we represented the placeholder word [P*] within the sentences using sks [30]. When extracting sentence features, the word embedding corresponding to the sks index is replaced with $P^{*}$.
@@ C Ablation Study on the Number of Inputs
Figure I: Ablation study on the number of input motions. Higher positions on the graph indicate better performance.
An ablation study on the number of inputs is conducted in the multiple input (MI) setting. As $|M_{i}|$ increases, all metrics demonstrate an increasing trend. This can be attributed to the higher likelihood of encountering motions that are more contextually aligned with the prompt as the diversity of input motions grows.
@@ D Persona Recognition Accuracy Details
For PerMo, PRA aims to classify 170 personas. To reduce the burden on a single classifier handling too many classes, we used nine separate classifiers, each responsible for a sub-parent category. Table D shows these sub-parent categories for PRA. For instance, the classifier for the "Age" category handles 20 personas, with each style containing five different personas.
Each classifier is a 2-block transformer-based model, pretrained separately for each sub-parent category. The classifiers are trained using ground truth motions in PerMo paired with persona labels. These classifiers evaluate the generated motions, and their classification accuracy determines the PRA. Note that all 170 personas are combined into a single dataset during training. For 100Style, Style Recognition Accuracy (SRA) is evaluated across 40 style categories, excluding content-oriented ones. Only one classifier is used for 100Style.
Table D: Sup-parent categories for measuring the PRA metric